### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. CHOIR’s workflow across different communication channels:
@@ 1. Introduction
To address this challenge, we present CHOIR (Chat-based Helper for Organizational Intelligence Repository), a chatbot-based organizational memory management system embedded in Slack. The primary goal of CHOIR is twofold: (1) to build and sustain a research lab’s organizational memory, and (2) to enable lab members to easily access knowledge from it. To that end, our approach embeds organizational memory directly within their communication platform—Slack—allowing students to ask questions to a chatbot and allowing professors to modify documents within the platform, leveraging knowledge that emerges from their conversations. To accomplish these goals, CHOIR integrates four interconnected interactions: Document-grounded Q&A, which retrieves knowledge with relevant references from a collection of documents when a user asks a question; Q&A Sharing for Follow-up Discussion, which allows users to share Q&A with peers to address unanswered questions or extend discussions; Knowledge Extraction from Conversation, which captures new knowledge that naturally emerges from Slack messages; and AI-assisted Document Update, which facilitates document maintenance by helping managers locate relevant documents and suggest edits. In this way, CHOIR supports an organizational memory practice that leverages lab members’ communication in Slack to retrieve and sustain knowledge, reducing lab directors’ burden of continuously updating documents while providing members with immediate and convenient access to organizational memory. While this approach can be potentially applicable to many different types of organizations, such as an online community and a company, we chose research labs as our target environment because they constantly experience member turnover, rely on regular communication, and maintain various types of documentation that require collaborative updates (Tuma et al., 2021).
@@ 2. Related Work
We explore prior research on organizational memory and knowledge management systems, examining how digital platforms have evolved to support knowledge preservation and sharing. We then investigate approaches to transforming conversational data into organizational knowledge, and finally review the emerging role of chatbots and large language models (LLMs) in facilitating knowledge work.
@@ 2.2. Leveraging Conversational Platforms for Knowledge Management
Despite these discoveries, converting conversations into persistent knowledge required addressing the challenge of preserving natural communication flow while enabling knowledge extraction. Early approaches focused on visualization and threading for improved retrieval (Donath et al., 1999; Kerr, 2003; Faridani et al., 2010), but required manual navigation. More sophisticated systems automated knowledge extraction through annotation (Ackerman et al., 2003; Fono and Baecker, 2006) and collaborative summarization (Zhang et al., 2017). Chat-based knowledge management systems achieved greater sophistication with Tilda, which automatically summarizes tagged conversations (Zhang and Cranshaw, 2018), and tools combining topic learning with personalized summaries (Tepper et al., 2018). QWiki demonstrated that Q&A and documentation approaches could be mutually reinforcing in MOOC environments (Setia et al., 2020). This trajectory confirms Ackerman’s (Ackerman et al., 2013) prediction that OM should move toward communication-centric, people-oriented approaches, while raising questions about scalability and sustainability.
@@ 3.1. Method
We recruited participants through a screening survey distributed to faculty and graduate students at research-oriented universities. We advertised the study on the author’s social media to recruit professors in other research-oriented universities. We recruited 15 participants: 11 faculty members (P1-P11) from various academic disciplines, primarily Computer Science (7) but also Industrial Engineering (2), Information Science (1), and Human Development and Family Science (1), and 4 graduate students (S1-S4) from different disciplines. We targeted professors who directed research labs with more than 4 Ph.D. students and had experience or interest in maintaining organizational memory. Faculty participants had been running their research labs for an average of 6.4 years (SD = 3.78, range: 4-13 years) and supervised an average of 5 Ph.D. students (SD = 1.67, range: 2-7 students). Some of the participants had a well-established organizational memory practice, whereas others relied on more informal organizational methods. This study was approved by our institution’s IRB([IRB-XXXXX]).
@@ 3.2. Result
“‘We have all kinds of gadgets, devices, and equipment. We have many robots, driving simulators, laptops, and eye trackers. And so they ask the same questions again and again.” (P6)
@@ 3.2.1. Challenges in Creating and Maintaining Lab Documentation
Design Goal 1 DG1: Facilitate the process of documenting organizational memory, leveraging the existing conversation among lab members.
@@ 3.2.2. Need for Effective Knowledge Retreival
Design Goal 2 DG2: Allow students to easily retrieve correct information and find the relevant document on demand via conversational interaction
@@ 3.2.3. Importance of Interpersonal Communication
Design Goal 3 DG3: Facilitate interpersonal communication and integrate the retrieval and update of organizational memory within the context of communication.
@@ 3.2.4. Documentation as a Collective Effort with Barriers
Design Goal 4 DG4: Support members’ participation in curating organizational memory and ensure accuracy.
These design goals guided the design and implementation of CHOIR, which we describe in the following section.
@@ 4. CHOIR: Chat-based Helper for Organizational Intelligence Repository
Figure 2. A questioner may ask CHOIR and receive a document-grounded answer, optionally share the Q&A exchange for broader discussion, and from these interactions, knowledge can be extracted and reviewed by the manager. Approved updates are written back to the organizational memory repository, which also supplies sources for Q&A.:
Figure 3. Document-grounded Q&A with CHOIR (DM example). Bob (questioner) asks CHOIR a question via direct message, CHOIR returns a document-grounded answer, and the thread is opened to view cited references.:
@@ 4.1. Document-grounded Q&A with CHOIR
Each Q&A exchange is posted with grounding and provenance so that the questioner can quickly verify sources. Figure 3 highlights the UI elements: the member’s question ( A), CHOIR’s answer ( B), an evidence panel listing the specific chunks used with anchored links ( C), and a share banner ( D) for further discussion. These references encourage the questioner to check out the original documents, thereby getting familiar with the organizational memory.
@@ 4.2. Q&A Sharing for Follow-up Discussion
Figure 4. Two share modals in CHOIR. Left: ”Share with Q&A” for posting a Q&A exchange to the designated channel. Right: ”Ask in Private” for sending the same exchange via DM; in this example, the questioner adds a follow-up, the manager is pre-selected in the People picker, and anonymous sharing is enabled.:
After CHOIR provides an answer, the questioner can share the Q&A exchange with others to facilitate communication among members (DG3). The share banner ( D in Figure 3) offers two options: posting to a public Q&A channel or sending via direct message to selected members. Questioners can configure additional options through one of two modal windows shown in Figure 4. The Your Comment field ( A in Figure 4) allows the questioner to add context or a follow-up prompt.
For the “Ask in Private” option, the question asker can choose lab members that they will share the question with using a People picker ( D in Figure 4), which is pre-populated with the manager by default but can be edited to include any Slack members (e.g., senior Ph.D. students). Privacy Options ( B) enable anonymous sharing, in which case CHOIR will replace the questioner’s name with “A team member” when the message is shared in a Q&A channel to stay anonymous. In addition, when anonymity is enabled in the private-sharing modal, the questioner is also excluded from the group chat conversation to preserve anonymity among recipients. The preview panel ( C) shows exactly how the question will be shared, either as a channel post or as an outgoing DM.
The shared Q&A exchange includes the question, CHOIR’s answer, the questioner’s comment, and profile image as shown in Figure 5 ( A), where other members can add follow-up comments ( B). When shared anonymously in a private DM, CHOIR forwards replies to the original questioner without revealing their identity. When such discussions lead to new knowledge, they provide the entry point for Knowledge Extraction from Conversation.
@@ 4.3. Knowledge Extraction from Conversation
Figure 5. Example of knowledge extraction after a Q&A is shared. The Q&A exchange has been posted to the designated channel (#qna) through the Share with Q&A modal. A colleague (Alice) adds a follow-up comment and then mentions CHOIR to request an update, prompting CHOIR to extract documentable knowledge from the conversation.:
CHOIR supports turning any Slack messages into reusable knowledge (DG1), allowing members to participate in organizational memory maintenance (DG4). Follow-up discussions after Q&A sharing often contain answers worth documenting. When any lab member sees new knowledge worth keeping, they can summon CHOIR by mentioning @CHOIR ( C in Figure 5). CHOIR then examines the conversation, identifies documentable content, and drafts a suggested update ( D). The requester can edit this draft and press Suggest Update to forward it to the manager ( E). CHOIR can be summoned anywhere in Slack—for example, if a professor shares a writing tip while giving feedback to a student, the professor can summon CHOIR to document that knowledge, as represented by the new start node in the bottom right corner of Figure 2.
@@ 4.4. AI-assisted Document Update
This feature of AI-assisted update supports DG1 and DG4 in multiple ways. Once any member suggests an update from the Knowledge Extraction feature by pressing the “Suggest Update” button ( E in Figure 5), CHOIR delivers it to the manager as a Slack DM message ( A in Figure 6) from CHOIR. The message presents the proposed update together with the contributor’s name and profile image, making the source and context explicit. At this stage, the manager can edit the suggestion, start the update process, or decline it, and these choices are visible not only in the DM but also as a message in the original channel so that other members are kept updated on how the update request is in progress. This final approval step, which the manager can only handle, is designed to help maintain the accuracy of documented information, which is relevant to DG4.
Once the manager starts the update process by pressing the “Start Update Process” button, CHOIR will send follow-up messages one step at a time, guiding the manager to access which file to update, and how. First, the manager can select a target file from the repository or create a new file if the knowledge does not fit into an existing one ( B in Figure 6). If not selected, CHOIR will use the most relevant file depending on the similarity of the suggested update and document content. CHOIR then shows which file is selected and the proposed change in a diff-style message with inline emphasis on additions and removals. For each suggestion, the manager may edit the content directly, apply the change to the document, skip it to show the next relevant document — as there can be multiple places to update in a document, or stop the review entirely ( C). If the knowledge does not fit into an existing section, the interface provides an option to create a new section through a dedicated modal ( D). Applied updates—whether to an existing file, a new file, or a new section—are written back to the repository, and CHOIR posts a message to the original channel acknowledging that the update has been finished.
This feature implements DG1 by automating key steps: recommending where updates are needed, proposing content, and committing changes. Managers can update documents directly in Slack without switching tools. CHOIR preserves managerial autonomy by allowing edits to proposed content or, for more complex tasks such as changing headings or reorganizing structure, enabling direct changes on GitHub via provided hyperlinks (“here” in D in Figure 6).
Figure 6. Manager-facing document update workflow in CHOIR. Left: a member’s suggestion (from Alice) arrives as a Slack DM with attribution and a link to the original discussion. Right: after starting the update process, the manager reviews the proposed change in context and decides how to incorporate it into the documentation.:
@@ 5. Field Study
To address RQ2, understanding how lab members engage with and experience a chatbot-driven organizational memory system, we conducted a month-long field deployment of CHOIR in the Slack workspaces of four university research labs, including the authors’ own.
@@ 5.1. Study Design
5.2. Participants and Recruitment
All participants provided informed consent for interviews, system usage monitoring, and interaction log collection. The study was approved by our institution’s IRB. As compensation, lab directors received \$40 and students received \$20 upon completing the interviews. One participating lab (L4) was led by the corresponding author, who served as the lab director. To mitigate any potential pressure that L4 participants might feel due to power dynamics or a desire to be favorable toward their peers, the director limited their role to answering questions and making changes to documents and did not otherwise intervene in the study procedures. Similarly, the first and second authors, who are members of L4, did not use CHOIR in this field deployment at all. All L4 participants voluntarily consented to the study independently through the standard procedure of advertisement, recruitment, and onboarding sessions, moderated by the first and second student authors. We further discuss potential limitations of this setup in the Discussion section.
Lab #. Department (Research Area) Lab Age Label (Gender) Status
L1 Computer Science 4 yrs D1 (M) Director
(Software Engineering) S1-1 (F) New student
@@ 5.2. Participants and Recruitment
L2 Computer Science 4 yrs D2 (F) Director
(Human-Computer Interaction) S2-1 (M) New student
@@
L3 Industrial Eng. 4 yrs D3 (F) Director
(Physical Ergonomics) S3-1 (M) New student
@@
L4 Computer Science 7 yrs D4 (M) Director (Author)
(Human-Computer Interaction) S4-1 (M) Existing student
@@
Table 1. Participants in the field study: each lab’s director and members, including newcomer status.:
@@ 5.3.2. Phase 2: Active Deployment
Following onboarding, CHOIR was deployed as a Slack App within each lab’s existing workspace. Participants were registered as authorized users with two distinct roles: lab directors served as managers with document approval access, while students operated as CHOIR users. Participant privacy was safeguarded through strict data collection protocols and automatic anonymization processes of personally identifiable information in compliance with FERPA requirements. When eligible messages were sent to external language models (OpenAI API), all usernames were replaced with pseudonyms, and responses were de-anonymized before delivery to users.

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791314isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. CHOIR’s workflow across different communication channels
This figure illustrates CHOIR’s workflow through three sequential interactions across different communication channels. Panel (1) shows a Questioner in a direct message with CHOIR asking about conference travel funding eligibility and receiving a document-grounded answer with source references, with options to share in the Q&A channel or ask privately. Panel (2) depicts a Peer in the public Q&A channel providing additional conference experience and suggesting a document update after CHOIR relays the question to peers or the manager. Panel (3) shows a Manager in a direct message with CHOIR receiving extracted knowledge and document update suggestions about conference funding support, with options to edit and apply changes. The bottom labels indicate how each role contributes to organizational memory: Questioner initiates inquiries, Peer provides collaborative input, and Manager approves document updates.
@@ 1. Introduction
To address this challenge, we present CHOIR (Chat-based Helper for Organizational Intelligence Repository), a chatbot-based organizational memory management system embedded in Slack. The primary goal of CHOIR is twofold: (1) to build and sustain a research lab’s organizational memory, and (2) to enable lab members to easily access knowledge from it. To that end, our approach embeds organizational memory directly within their communication platform, such as Slack, allowing students to ask questions to a chatbot and allowing professors to modify documents within the platform. Specifically, CHOIR integrates four interconnected interactions: Document-grounded Q&A, which retrieves knowledge with relevant references from a collection of documents when a user asks a question; Q&A Sharing for Follow-up Discussion, which allows users to share Q&A with peers to address unanswered questions or extend discussions; Knowledge Extraction from Conversation, which captures new knowledge that naturally emerges from Slack messages; and AI-assisted Document Update, which facilitates document maintenance by helping managers locate relevant documents and suggest edits. In this way, CHOIR supports an organizational memory practice that leverages lab members’ communication in Slack to retrieve and sustain knowledge, reducing lab directors’ burden of continuously updating documents while providing members with immediate and convenient access to organizational memory. While this approach can be potentially applicable to many different types of organizations, such as an online community and a company, we chose research labs as our target environment because they constantly experience member turnover, rely on regular communication, and maintain various types of documentation that require collaborative updates (Tuma et al., 2021).
@@ 2. Related Work
While prior work on personal knowledge management (Dumais et al., 2003; Gemmell et al., 2006; Capra and Perez-Quinones, 2005) focused on search-based retrieval across collected sources, an organization’s knowledge management is challenging due to its scale and group dynamics. We explore prior research on organizational memory and knowledge management systems, examining how digital platforms have evolved to support knowledge preservation and sharing. We then investigate approaches to transforming conversational data into organizational knowledge, and finally review the emerging role of chatbots and large language models (LLMs) in facilitating knowledge work.
@@ 2.2. Leveraging Conversational Platforms for Knowledge Management
Despite these discoveries, converting conversations into persistent knowledge required addressing the challenge of preserving natural communication flow while enabling knowledge extraction. Early approaches focused on visualization and threading for improved retrieval (Donath et al., 1999; Kerr, 2003; Faridani et al., 2010), but required manual navigation. More sophisticated systems automated knowledge extraction through annotation (Ackerman et al., 2003; Fono and Baecker, 2006) and collaborative summarization (Zhang et al., 2017). Chat-based knowledge management systems achieved greater sophistication with Tilda, which automatically summarizes tagged conversations (Zhang and Cranshaw, 2018), and tools combining topic learning with personalized summaries (Tepper et al., 2018). Social Q&A systems demonstrated how questions can surface knowledge gaps and how answers gradually accumulate into shared resources (Richardson and White, 2011). QWiki demonstrated that Q&A and documentation approaches could be mutually reinforcing in MOOC environments (Setia et al., 2020). This trajectory confirms Ackerman’s (Ackerman et al., 2013) prediction that OM should move toward communication-centric, people-oriented approaches, while raising questions about scalability and sustainability.
@@ 3.1. Method
We recruited participants through a screening survey distributed to faculty and graduate students at research-oriented universities. We advertised the study on the author’s social media to recruit professors in other research-oriented universities. We recruited 15 participants: 11 faculty members (P1-P11) from various academic disciplines, primarily Computer Science (7) but also Industrial Engineering (2), Information Science (1), and Human Development and Family Science (1), and 4 graduate students (S1-S4) from different disciplines. We targeted professors who directed research labs with more than 4 Ph.D. students and had experience or interest in maintaining organizational memory. Faculty participants had been running their research labs for an average of 6.4 years (SD = 3.78, range: 4-13 years) and supervised an average of 5 Ph.D. students (SD = 1.67, range: 2-7 students). Some of the participants had a well-established organizational memory practice, whereas others relied on more informal organizational methods. This study was approved by the Virginia Tech Institutional Review Board (IRB 25-247).
@@ 3.2. Result
“We have all kinds of gadgets, devices, and equipment. We have many robots, driving simulators, laptops, and eye trackers. And so they ask the same questions again and again.” (P6)
@@ 3.2.1. Challenges in Creating and Maintaining Lab Documentation
Design Goal 1 (DG1): Facilitate the process of documenting organizational memory, leveraging the existing conversation among lab members.
@@ 3.2.2. Need for Effective Knowledge Retreival
Design Goal 2 (DG2): Allow students to easily retrieve correct information and find the relevant document on demand via conversational interaction.
@@ 3.2.3. Importance of Interpersonal Communication
Design Goal 3 (DG3): Facilitate interpersonal communication and integrate the retrieval and update of organizational memory within the context of communication.
@@ 3.2.4. Documentation as a Collective Effort with Barriers
Design Goal 4 (DG4): Support members’ participation in curating organizational memory and ensure accuracy.
Figure 2. Overview of CHOIR’s workflow. A questioner may ask CHOIR and receive a document-grounded answer, optionally share the Q&A exchange for broader discussion, and from these interactions, knowledge can be extracted and reviewed by the manager. Approved updates are written back to the organizational memory repository, which also supplies sources for Q&A.
Flow diagram with four main components: Document-grounded Q&A with CHOIR (actor: questioner asking a question); Q&A Sharing for Follow-up Discussion (actors: questioner sharing the Q&A exchange, peers and the manager contributing insights); Knowledge Extraction from Conversation (actors: peers or the manager suggesting CHOIR save knowledge, or directly providing new knowledge); and AI-assisted Document Update (actor: the manager reviewing and applying changes). The organizational memory repository is placed centrally, connected by READ arrows for Q&A retrieval and UPDATE arrows for approved changes. Each decision node is controlled by human actors—questioners, peers, or the manager—and branches exist where no updates are made if actors choose not to proceed.
Figure 3. Document-grounded Q&A with CHOIR (DM example). Bob (questioner) asks CHOIR a question via direct message, CHOIR returns a document-grounded answer, and the thread is opened to view cited references.
Slack UI in DM mode, with four numbered callouts: (1) member’s question submitted via direct message; (2) CHOIR’s grounded answer in the main pane; (3) evidence panel in the thread listing References 1–3 with quoted snippets and source links; (4) share banner, visible only to the questioner, offering two buttons: ”Ask the Q&A Channel” and ”Ask in Private.” This illustrates the DM case, though questions can also be submitted by mentioning CHOIR in a channel.
@@ 4. CHOIR: Chat-based Helper for Organizational Intelligence Repository
Figure 4. Two share modals in CHOIR. Left: ”Share with Q&A” for posting a Q&A exchange to the designated channel. Right: ”Ask in Private” for sending the same exchange via DM; in this example, the questioner adds a follow-up, the manager is pre-selected in the People picker, and anonymous sharing is enabled.
Side-by-side Slack modals. Left (”Share with Q&A”): (1) optional comment field; (2) anonymous sharing option; (4) preview of the channel post including the Q&A exchange. Primary action: ”Post to Channel.” Right (”Ask in Private”): (1) comment field containing an additional follow-up; (3) People picker pre-populated with the manager (editable); (2) anonymous option checked (excludes the questioner from the DM and shows ”A team member”); (4) preview of the outgoing DM. This figure illustrates a DM-started case; questions can also be asked by mentioning CHOIR in a channel.
Figure 5. Example of knowledge extraction after a Q&A is shared. The Q&A exchange has been posted to the designated channel through the Share with Q&A modal. A peer (Alice) adds a follow-up comment and then mentions CHOIR to request an update, prompting CHOIR to extract documentable knowledge from the conversation.
Left: the shared Q&A appears in the qna channel, showing the question, CHOIR’s answer, the questioner’s comment, and the profile image (or ”A member” if anonymous). Below, Alice posts a follow-up comment. Right: Alice mentions CHOIR to suggest an update, and CHOIR responds with a draft ”Suggested Update” block that can be edited and submitted to the manager.
Figure 6. Manager-facing document update workflow in CHOIR. Left: a member’s suggestion (from Alice) arrives as a Slack DM with attribution and a link to the original discussion. Right: after starting the update process, the manager reviews the proposed change in context and decides how to incorporate it into the documentation.
Two-panel screenshot. Left: proposal intake card with author attribution, suggestion text, and controls (Edit Suggestion, Start Update Process, Decline), followed by a file picker and Create New File. Right: diff-like suggestion view with bold and strikethrough to indicate edits within a named file and section; action buttons include Edit, Apply, Skip, Stop, plus a Create New Section option and a link to edit in GitHub.
@@ 4.1. Document-grounded Q&A with CHOIR
[…]
@@ 4.2. Q&A Sharing for Follow-up Discussion
[…]
@@ 4.3. Knowledge Extraction from Conversation
[…]
@@ 4.4. AI-assisted Document Update
[…]
@@ 5. Field Study
To understand how lab members engage with and experience a chatbot-driven organizational memory system (RQ2), we conducted a month-long field deployment of CHOIR in the Slack workspaces of four university research labs, including the authors’ own.
@@ 5.1. Study Design
Table 1. Participants in the field study: each lab’s director and members, including newcomer status. Lab age is shown in parentheses below each lab number.
A table showing 21 participants across 4 research labs (L1-L4) with 4 columns: Lab number with age, Department and research area, Participant label with gender, and Status. L1 is a 4-year-old Software Engineering lab with 1 director (D1, male) and 4 members (1 new female member, 3 existing members). L2 is a 4-year-old HCI lab with 1 director (D2, female) and 5 members (2 new male members, 3 existing members). L3 is a 4-year-old Physical Ergonomics lab with 1 director (D3, female) and 5 members (2 new members, 3 existing members). L4 is a 7-year-old HCI lab with 1 director who is also an author (D4, male) and 4 existing members. The table shows a mix of new and existing members across all labs, with varied gender representation.
Lab [1pt](years) Department [1pt](Research Area) Label [1pt](Gender) Status
L1 Computer Science D1 (M) Director
(4 yrs) (Software Engineering) S1-1 (F) New student
@@
L2 Computer Science D2 (F) Director
(4 yrs) (HCI) S2-1 (M) New student
@@
L3 Industrial Engineering D3 (F) Director
(4 yrs) (Physical Ergonomics) S3-1 (M) New student
@@
L4 Computer Science D4 (M) Director (Author)
(7 yrs) (HCI) S4-1 (M) Existing student
@@
5.2. Participants and Recruitment
All participants provided informed consent for interviews, system usage monitoring, and interaction log collection. The study received IRB approval (IRB 25-247). As compensation, lab directors received \$40 and students received \$20 upon completing the interviews. One participating lab (L4) was led by the corresponding author, who served as the lab director. In addition to easing recruitment, the study was motivated by the director’s experience with manually curating organizational memory and observing the limited use of these materials in practice. To mitigate any potential pressure that L4 participants might feel due to power dynamics or a desire to be favorable toward their peers, the director limited their role to answering questions and making changes to documents and did not otherwise intervene in the study procedures. Similarly, the first and second authors, who are members of L4, did not use CHOIR in this field deployment at all. All L4 participants voluntarily consented to the study independently through the standard procedure of advertisement, recruitment, and onboarding sessions, moderated by the first and second authors.
@@ 5.3.2. Phase 2: Active Deployment
Following onboarding, CHOIR was deployed as a Slack App within each lab’s existing workspace. Participants were registered as authorized users with two distinct roles: lab directors served as managers with document approval access, while students operated as CHOIR users. No specific usage requirements were imposed on participants, while awareness of CHOIR’s availability was maintained by having the research team periodically ask questions to CHOIR in shared channels. Participant privacy was safeguarded through strict data collection protocols and automatic anonymization processes of personally identifiable information in compliance with FERPA requirements. When eligible messages were sent to external language models (OpenAI API), all usernames were replaced with pseudonyms, and responses were de-anonymized before delivery to users.